# VersiCode: Towards Version-controllable Code Generation

### v1

@@ 3.2 Results and Analysis
The programming knowledge of LLMs, particularly regarding version-specific information, is surprisingly outdated. In Figure 5 we present the Pass@1 performance for token-level code completion, grouped by year, covering 2015-2023. Additionally, we show the histogram of data distribution for each year. To ensure precise timestamps and minimize noise, we only used instances collected from library source code. As shown in Figure 5-a we can observe a general trend: the later the time, the worse the models’ performance. This is counter-intuitive compared to temporal knowledge question answering [52], where performance initially increases before declining. We further filtered for “deprecation” (Figure 5-b) and “addition” (Figure 5-c) to identify version-sensitive cases. While the sparsity of data decreases confidence in results, in both cases, we can observe a consistent decreasing trend over time. This suggests that LLMs have outdated programming knowledge, highlighting the need for rapid adaptation to newer libraries and APIs.
@@ 4 Discussion
Can we address version-controllable code generation with retrieval-augmented generation (RAG)? RAG approaches typically involve two crucial components: retrieval and in-context generation [11]. From the retrieval perspective: (1) The difficulty in disambiguating version-related queries, as embeddings for versions like "torch 2.1.3" and "torch 1.3.2" can be very similar [39]. This similarity makes it hard for models to differentiate between specific features and capabilities associated with each version. ( 2) Version information of code snippets is rarely explicitly mentioned within the code itself and may instead appear in separate configuration files like requirements.txt. This separation necessitates a more sophisticated retrieval approach, where the model must integrate information from multiple sources to accurately understand the version dependencies. However, from the perspective of in-context generation: Table 3 shows that even non-matching version contexts (i.e., code editing) can help smaller models generate grammatically correct code. The observation suggests potential for dedicated RAG approaches [18], though the benefits are limited and retrieval noise may reduce effectiveness.
@@
How can we enhance pre-training for new code-LLMs? Figure 5 demonstrates a notable decline in the performance of all models over time. This deterioration is likely attributable to two primary factors: (1) the use of outdated pre-training data, which causes older versions of code to predominate the training set, and (2) the backward compatibility of APIs, which results in a higher prevalence of use cases and examples pertaining to older versions of these APIs [22]. To mitigate this issue and improve the models’ capabilities with newer libraries, we suggest increasing the representation of new-version codebases within the training data. This adjustment aims to enhance the models’ proficiency in utilizing contemporary libraries effectively [52, 38].
@@
How can we address the challenge of evolving libraries in LLMs? Generating block-level or repository-level code [30] requires LLMs to understand user demands and library dependencies. Addressing this challenge involves continually training the model with new libraries using continual learning techniques [18]. These techniques enable the model to adapt to changing libraries without forgetting previously learned information. Examples include memory-based methods and various continual learning strategies [44, 46, 43]. Additionally, developing benchmark datasets that are continuously and automatically curated and maintained is crucial for evaluating the performance of models with new libraries [16]. Enriching the taxonomy [19] and maintaining datasets for evolving libraries [22] is also vital [19]. Multi-agent systems can be employed for this purpose. Aligning development and evaluation efforts will enhance the ability of LLMs in code understanding and generation capabilities, ensuring they remain effective as libraries evolve.
@@ 5 Related Work
Code Generation Models: Recent advancements in code language models [13, 7, 5, 37, 40], driven by sophisticated NLP techniques [18] and extensive code repositories [15], have resulted in substantial breakthroughs. Transformer-based large language models [32, 37, 13, 28, 5, 12, 24] have demonstrated exceptional capabilities in generating syntactically correct and semantically meaningful code from natural language descriptions. Additionally, research efforts that integrate multi-modal data [34, 36, 33], including both code and accompanying documentation [15], have significantly improved model accuracy. While in real-world software engineering,
Code Generation Datasets: The code generation [18, 40, 31] includes tasks for both code completion and code editing, ensuring comprehensive coverage of programming scenarios. Code completion [48, 49, 8, 6, 4, 14, 29, 25, 9, 27, 21, 50, 10, 53] is the task of predicting subsequent code tokens based on the given context, benefits from datasets, which provide extensive code repositories from various programming languages. These datasets enable models to learn syntactic and semantic patterns [19]. Code editing [20, 26, 55, 54, 15, 47, 2, 19, 51, 41] involves automatically generating changes to existing code, such as bug fixes or refactoring. Datasets like EvalGPTFix [51] and DebugBench [41], which focus on bug fixing and code refinement tasks, are instrumental in this area. To our knowledge, given the necessity and challenges in library evolution [18], refer to the detailed comparison in Table 1 and Appendix B.2, the proposed dataset VersiCode is the first large-scale code generation dataset, covering both code completion and code editing.
@@ 6 Conclusion
In this paper, we introduce the VersiCode dataset, a large-scale dataset of version-controllable Python code. VersiCode encompasses 300 libraries across more than 2,000 versions spanning 9 years, with a total of more than 100,000 data points. In conjunction with VersiCode, we designed two novel tasks aimed to test the capabilities of large language models in handling version-controlled code completion and editing tasks. The comprehensive experiments reveal that current models struggle significantly with these complex tasks. While token-level tasks were less challenging, block-level tasks highlighted substantial weaknesses in handling detailed coding structures. These findings highlight the need for ongoing research to better manage the dynamic nature of source code and its evolution. We plan to expand the VersiCode dataset to include more programming languages and detailed version histories. We believe version-controllable code generation is crucial and practical, and the VersiCode dataset is valuable for advancing LLMs toward deployment in professional software development.
7 Limitation
While the construction of VersiCode involved LLM-generated annotations and a partially automated pipeline, it still requires a certain amount of human annotation work for quality control. Fully automating the dataset curation and maintenance process [45] requires further development. This capability is especially important given the potential risk of dataset contamination [23] and the ever-evolving nature of libraries and APIs [22]. Another limitation of the current VersiCode dataset is its handling of multiple programming languages. While we have conducted annotations on Java, JavaScript, and C# from Stack Overflow, as shown in Appendix B.1 and  E.2, the data size for these languages is not comparable to Python. More effort, similar to that invested in constructing the Python version of VersiCode, is required for other languages. The final limitation of the current VersiCode dataset is the simplicity of the editing scheme [19], improving which may bridge the gap between code generation and real-world software engineering tasks.
Appendix A Annotation Rules
As depicted in Table 5, we have detailed the annotation and filtering rules employed during the preprocessing of raw data.
Table 5: Annotation and filtering rules.
@@ Appendix A Annotation Rules
Ranked Libraries ->StackOverflow Q&A
@@
Ranked Libraries ->Library Source Code Based on the ranked libraries, parse the source code of these libraries to find functions related to version changes.
Ranked Libraries ->Downstream Application
@@
Appendix B Related Dataset
B.1 Detailed Data Statistics of VersiCode
As shown in Table 6, we have outlined the complete version of VersiCode in the table, which furnishes human-labeled data for three additional languages: C#, Java, and JavaScript.

### v2

@@ 3.2 Results and Analysis
The programming knowledge of LLMs, particularly regarding version-specific information, is surprisingly outdated. Figure 4 presents the EM@1 performance for token-level code completion, grouped by year from 2015 to 2023, along with a histogram showing the data distribution for each year. To ensure precise timestamps and minimize noise, we only used instances collected from library source code. As shown in Figure 4-a, there is a clear trend: model performance declines as the release time becomes more recent. This is counter-intuitive compared to temporal knowledge question answering (Zhao et al., 2024), where performance initially increases before declining. We further filtered for “deprecation” (Figure 4-b) and “addition” (Figure 4-c) to identify version-sensitive cases. Although data sparsity reduces confidence in the results, both cases show a clear downward trend over time This suggests that LLMs have outdated programming knowledge, highlighting the need for rapid adaptation to newer libraries and APIs.
4 From Token-level to Line- and Block-level Completion
When utilizing third-party code library APIs, LLMs should handle not only API name generation but also parameter preparation and contextual code integration. In this section, we extend the task to line-level (completing a single line) and block-level (completing multiple lines) code generation. This expanded scope presents new challenges for both the model’s capabilities and the evaluation methodologies. (1) How does increasing complexity in line- or block-level code completion affect the LLMs to handle API usage and parameters? (2) How does having more context (like import statements and specified library version) improve the accuracy of line- and block-level code generation? (3) Which evaluation metrics best capture the accuracy of line- and block-level code generation, and which is most reliable?
4.1 Experiment Setup
Models: We selected GPT-4o, GPT-3.5, and LLaMA3 70B, the three models that perform best on token-level code completion, to conduct experiments on line-level or block-level code completion.
Data Preparation: We sample a subset from VersiCode for dynamic code analysis with executable test cases from library source code, focusing on code snippets with complete context (e.g., import statements). GPT-4 was used to refactor the snippets into task functions, followed by test case generation and validation in a version-specific environment. All of the test cases have been manually verified to ensure their correctness. The code completion tasks are categorized into token, line, and block levels. The test cases include return type, normal input, boundary values, and functionality checks (see Appendix A.3 for details).
Figure 5: The process of executable code assessment, which includes data refactoring, test case generation, and validation. Starting from code snippets collected from real code involving specific API calls for a given library version, GPT-4 is employed to refactor the code into a task function. The large language model is then prompted to generate test cases from various perspectives (See Appendix F for a running example of instances and test cases.) Each generated test case is verified by experts, and the correctness is ensured by running the code in a specified environment. If issues arise, they are corrected through multiple iterations with GPT-4.
Metrics: We use the following evaluation metrics for each task granularity: (1) Pass@$k$ for token-level generation (Chen et al., 2021): For this metric, we generate $n\geq k$ samples per instance (with $n=6$ and $k=1$ to compare different metrics). We count the number of correct samples $c\leq n$ judged by executable testing. (2) Identifier Sequence Match (ISM@$k$) and Prefix Match (PM@$k$) for line-level generation (Agrawal et al., 2023): These metrics measure how closely the generated sequences match the ground truth. For block-level generation, we adopt the average performance over lines. Following the setup in Agrawal et al. (Agrawal et al., 2023), we generate $n=6$ independent samples per instance. (3) Exact Match (EM@$k$): We use regular expression matching to determine whether the specified API is used in the code generated by the model and the formula for calculating the EM@k score is the same as the formula for calculating the Pass@k score ($n=6$ and $k=1$). (4) Critical Diff Check (CDC@$k$): Unlike traditional code similarity calculations, CDC focuses on the differences between the code generated by the model and the reference answer. CDC extends the EM metric by adding four additional rules: checking whether the generated code is syntactically valid; identifying the line in the generated code where the specified API is used and determining if the number of parameters in the function call is the same; if the answer uses a with statement, checking whether the generated code also uses a with statement; and if the answer uses keyword arguments, verifying whether the generated code uses the same keyword arguments. Please refer to Appendix E for detailed examples, effectiveness analysis, and ablation study, conducted to validate CDC.
4.2 Results and Analysis
Figure 6: The Pass@1 performance of different models across various granularities and test case types.
Table 1: The performance of different models across various granularities (Token, Line, Block). Pass@1 refers to dynamic analysis metrics, while green-colored metrics (EM, ISM, PM) correspond to static analysis based on string matching. The blue-colored metric (CDC) represents a newly proposed metric. The configurations labeled as “w/o version” indicate that the prompt does not specify the version of the third-party code libraries, while “w/o import” refers to prompts where the provided code context lacks import statements, meaning the model must generate code based entirely on user intent. The Pearson correlation coefficient is computed for each metric’s results against Pass@1 within each granularity.
Less context leads to more errors in code generation. When models have more context, like import statements, their performance improves significantly. For example, as shown in Table 1, GPT-4o at the token level achieves a Pass@1 score of 65.97 with imports, but this drops to 44.54 without imports. This pattern is consistent across all models and granularity levels (i.e., token, line, and block), as shown in Figure 6. When models lack important context, such as external libraries or other dependencies, they struggle to generate accurate code, which leads to more errors. So, giving models more information upfront is crucial for better results.
Models show limited sensitivity to version-specified instructions. As shown in Table 1, at the token level, models like GPT-4o perform slightly better when provided with version information (52.80 with version v.s. 49.72 without version). However, this advantage diminishes at the line and block levels, where the results become inconsistent. This suggests that while version details can be helpful for short code snippets, they don’t significantly impact the model’s performance for more extended or complex code. This likely indicates that models are not trained to prioritize or heavily rely on version-specific instructions.
The CDC@1 metric closely aligns with Pass@1 scores, making it a strong proxy for dynamic code analysis. As shown in Table 1, at the block level, the Pearson Correlation Coefficient (PCC) between CDC@1 and Pass@1 is 0.9995, indicating a strong correlation. Even though EM@1 has a high correlation with Pass@1 at the token level (PCC = 0.9995), EM@1 becomes less aligned at the block level (PCC = 0.8974). Additionally, the absolute differences between CDC@1 and Pass@1 values are generally smaller compared to other static metrics like EM@1, making CDC a potentially more reliable alternative for assessing code generation accuracy.
5 From Code Completion to Code Migration
In addition to generating code for specific third-party library versions, another common challenge is maintaining user projects when these libraries are upgraded or rolled back. We address version-aware code migration by exploring three key questions: (1) How well can LLMs handle migrating code across different versions, compared to generating code for a specific version? (2) What impact do major and minor version changes in third-party libraries have on code migration? (3) How do forward migrations (from older to newer versions) compare to reverse migrations (from newer to older versions) in terms of trends and challenges?
5.1 Experiment Setup
Models: Based on the token-level code completion experimental results in Section 3, we selected the most outstanding performers from each model series for the experiments in this section.
Data Preparation: For code migration, we utilize a subset of VersiCode, in which instances are constructed based on differences between source and target code versions, covering both updates to newer versions and downgrades. Versions were categorized by patterns (e.g., major vs. minor) to capture different migration scenarios. (Detailed in Appendix A.3)
Metrics: Code migration is similar to block-level tasks in code completion. We use the same evaluation metric as for block-level: CDC@$k$ ($n=6$, $k\in\{1,3\}$).
5.2 Results and Analysis
Table 2: The performance of various models in different code migration scenarios. The arrow $\mapsto$” indicates the direction of migration, where ”Major” corresponds to major version changes (e.g., Torch 2.0.0), and ”Minor” corresponds to minor version changes (e.g., Torch 2.1.3). The ”Old $\mapsto$ New” scenario simulates upgrading from an old version to a new version, while ”New $\mapsto$ Old” represents the maintenance of historical code. The performance of different models in these scenarios is measured using the CDC metrics (CDC@1 and CDC@3), reflecting their adaptability to various code migration tasks.
Model performance across version migrations. Different models display varying degrees of adaptability when transitioning between major and minor software versions, with some showing exceptional robustness in Table 2. The table categorizes version migrations into four types: Major-to-Major, Major-to-Minor, Minor-to-Major, and Minor-to-Minor. Notably, models like GPT-4o excel in major-to-major migrations, suggesting superior handling of significant changes, whereas transitions involving minor versions tend to show moderate performance. This variability underscores the models’ design focus, whether on broad adaptability or specialized functionality.
Adaptability in code migration based on release timing. Backward and forward compatibility testing reveals a spectrum of model resilience under different temporal migration scenarios. The evaluation is split into two releasing time directions: Old-to-New and New-to-Old, shown in Table 2. Generally, models perform better when adapting to newer versions from older ones, with GPT-4o standing out for its high scores in both directions. However, the drop in performance when handling older versions after training on newer releases highlights challenges in maintaining backward compatibility, a critical aspect for long-term usability and integration stability in evolving tech environments.
The context code in another version is still helpful, but its benefits are limited. The comparison between block-level code completion and block-level code migration is shown in Table 2 and Table 1. There is a significant improvement across most models, except for LLaMA3-70B and GPT-4o, detailed in Appendix D.3. When provided with code in another version as context (i.e. in the code migration task), these models can generate correct code with a much higher success rate. However, a bottleneck is evident in LLaMA3-70B and GPT-4o, where the code context hinders their performance than code completion.
@@ 6 Discussion
How can we enhance pre-training for new code-LLMs? Figure 4 demonstrates a notable decline in the performance of all models over time. This deterioration is likely attributable to two primary factors: (1) the use of outdated pre-training data, which causes older versions of code to predominate the training set, and (2) the backward compatibility of APIs, which results in a higher prevalence of use cases and examples about older versions of these APIs (Lamothe et al., 2022). To mitigate this issue and improve the models’ capabilities with newer libraries, we suggest increasing the representation of new-version codebases within the training data. This adjustment aims to enhance the proficiency in utilizing contemporary libraries effectively (Zhao et al., 2024; Shao et al., 2024). Besides, based on the results in Section 4.2, current LLMs show limited use of version information in code generation. To address this, we propose enhancing pre-training by incorporating version-tagged code samples and metadata to help models better differentiate between API versions.
@@
How can we address the challenge of evolving libraries in LLMs? Generating block-level or repository-level code (Luo et al., 2024a) requires LLMs to understand user demands and library dependencies. Addressing this challenge involves continually training the model with new libraries using continual learning techniques (Jiang et al., 2024). These techniques enable the model to adapt to changing libraries without forgetting previously learned information. Examples include memory-based methods and various continual learning strategies (Wu et al., 2024; Yadav et al., 2023; Wu et al., 2022). Additionally, developing benchmark datasets that are continuously and automatically curated and maintained is crucial for evaluating the performance of models with new libraries (Jang et al., 2022). Enriching the taxonomy (Jiao et al., 2023) and maintaining datasets for evolving libraries (Lamothe et al., 2022) is also vital (Jiao et al., 2023). Multi-agent systems can be employed for this purpose. Aligning development and evaluation efforts will enhance the ability of LLMs in code understanding and generation capabilities, to remain effective as libraries evolve.
Can we address version-controllable code generation with retrieval-augmented generation? Retrieval-augmented generation (RAG) approaches typically involve two crucial components: retrieval and in-context generation (Gao et al., 2023). The following challenges need to be addressed in order for RAG to be effectively applied to this problem. From the retrieval perspective: (1) It may be difficult to disambiguate version-related queries, as embeddings for version strings like “torch 2.1.3” and “torch 1.3.2” can be very similar (Singh & Strouse, 2024). This similarity makes it hard for retrievers to differentiate between specific features and capabilities associated with each version. ( 2) Version information of code snippets is rarely explicitly mentioned within the code itself and may instead appear in separate configuration files like “requirements.txt”. This separation necessitates a more sophisticated retrieval approach, where the model must integrate information from multiple sources to accurately understand version dependencies. From the perspective of in-context generation: Table 9 shows that even non-matching version contexts (i.e., code migration) can help smaller models generate grammatically correct code. This observation suggests potential for dedicated RAG approaches (Jiang et al., 2024), though the benefits are limited and retrieval noise may reduce effectiveness.
What are the effective methods for evaluating the capabilities of LLMs in generating version-controllable code? Both static analysis (Agrawal et al., 2023), which reviews code without executing it, and dynamic analysis (Zhuo et al., 2024), which tests the code by running it, are vital for software development. However, evaluating LLMs for version-controllable code generation presents unique challenges. (1) Dynamic analysis is complicated by API calls that rely on specific code contexts, making it difficult and costly to create standalone tests (Zhuo et al., 2024). Additionally, using LLM-generated code as test cases introduces further complexity in managing test quality. Especially, VersiCode, which includes 300 packages and over 2,000 versions in the raw dataset, requires detailed setups for each testing environment and managing various dependencies, complicating the practical deployment of solutions. (2) Meanwhile, static analysis uses metrics like ISM (Agrawal et al., 2023) and PM (Agrawal et al., 2023) for broad coverage but may miss critical details such as indentation and parameter positioning in API-related code, refer to Table 1 and Appendix E. These omissions suggest that traditional static metrics are not entirely suitable for assessing version-controllable code generation. Evaluating the effectiveness of these metrics is crucial. Our study initiates the exploration of more reliable methods; however, extensive research, including approaches like code slicing (Du et al., 2024), is essential to advance our evaluation techniques.
@@ 7 Related Work
Code Generation Models: Recent advancements in code language models (Guo et al., 2024; CodeGemma Team et al., 2024; Bai et al., 2023; Rozière et al., 2023; Sun et al., 2024), driven by sophisticated NLP techniques (Jiang et al., 2024) and extensive code repositories (Hu et al., 2023), have resulted in substantial breakthroughs. Transformer-based large language models (Luo et al., 2024c; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Bai et al., 2023; Gunasekar et al., 2023; Li et al., 2023) have demonstrated exceptional capabilities in generating syntactically correct and semantically meaningful code from natural language descriptions. Additionally, research efforts that integrate multi-modal data (OpenAI, 2023b; 2024; Meta LlaMa team, 2024), including both code and accompanying documentation (Hu et al., 2023), have significantly improved model accuracy. While in real-world software engineering,
Code Generation Datasets: The code generation (Jiang et al., 2024; Sun et al., 2024; Luo et al., 2024b) includes tasks for both code completion and code editing, ensuring comprehensive coverage of programming scenarios. Code completion (Yao et al., 2018; Yin et al., 2018; Feng et al., 2020; Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Lu et al., 2021; Li et al., 2022; Fried et al., 2023; Liu et al., 2023; Lai et al., 2023; Yu et al., 2024; Fu et al., 2023; Zheng et al., 2023) is the task of predicting subsequent code tokens based on the given context, benefits from datasets, which provide extensive code repositories from various programming languages. These datasets enable models to learn syntactic and semantic patterns (Jiao et al., 2023). Code editing (Just et al., 2014; Lin et al., 2017; Zhu et al., 2022b; a; Hu et al., 2023; Yan et al., 2023; Ahmad et al., 2023; Jiao et al., 2023; Zhang et al., 2023; Tian et al., 2024) involves automatically generating changes to existing code, such as bug fixes or refactoring. Datasets like EvalGPTFix (Zhang et al., 2023) and DebugBench (Tian et al., 2024), which focus on bug fixing and code refinement tasks, are instrumental in this area. To our knowledge, given the necessity and challenges in library evolution (Jiang et al., 2024), refer to the detailed comparison in Table 6 and Appendix C, the proposed dataset VersiCode is the first large-scale code generation dataset, covering both code completion and code editing. Refer to Appendix C for a comprehensive comparison among datasets.
Third-party Library Evolution: Third-party library code is continually updated due to bug fixes, code refactoring, and the addition of new features, making it a significant research topic in software engineering (Zhang et al., 2020; 2021; Dilhara et al., 2021; Liu et al., 2021; Wang et al., 2020; Vadlamani et al., 2021; Haryono et al., 2021). Studies by Zhang et al. (2020) show that Python APIs often evolve by adding, deleting, or modifying parameters. Further research by Zhang et al. (2021) notes frequent API changes, including parameter updates. Dilhara et al. (2021) reveal that developers adjust their use of machine learning libraries in response to updates, while Liu et al. (2021) and Dig & Johnson (2006) find that undocumented changes in Android and Java can cause errors. Research on API deprecation highlights issues with documentation and the quality of suggested alternatives (Wang et al., 2020; Vadlamani et al., 2021; Haryono et al., 2021; Brito et al., 2018), showing that improvement in library evolution does not necessarily translate to better suggestions for deprecated APIs. VersiCode, unlike traditional software engineering research, studies API version evolution from an LLM perspective, exploring its impact on model training, code generation, and evaluation.
@@ 8 Conclusion
In conclusion, our research underscores the need for updated benchmarks that capture the dynamic nature of software development, better assessing the capabilities of LLMs in code generation. By introducing the VersiCode dataset, we provide a realistic testing ground that reveals significant limitations in current models, like GPT-4o and LLaMA3, when handling version-specific code. Our findings advocate for continuous model improvements and the adoption of our new metric, i.e., critical diff check, which more accurately evaluates model performance against real-world challenges. This work not only introduces valuable tools but also sets a direction for future enhancements in AI-driven code generation, ensuring LLMs remain effective and relevant in professional settings. For future research, we will investigate a solution for version-controllable code generation based on the insights from this paper, including approaches like continual learning, memory-enhanced methods, or retrieval-based methods. Additionally, we plan to develop a live version of VersiCode, which will continuously incorporate new libraries and downstream use cases.
Appendix A Dataset Construction
VersiCode is a large-scale code generation benchmark dataset focusing on evolving library dependencies. We propose two tasks to simulate real-world applications: version-specific code completion and version-aware code migration, incorporating version information into code generation constraints. First, we discuss data curation, and preprocessing of noisy code snippets and FAQs into organized metadata. Based on the metadata, we describe the task design and quality control process. We then address tagging API lifespan features per library version. Finally, we provide data statistics for VersiCode and discuss future dataset extensions.
A.1 Dataset Curation and Collection
As shown in Figure 7, we first collected permissively licensed Python repositories from GitHub that serve as the source code for Python libraries. These repositories are ranked by their popularity (as indicated by their collected stars). Using the list of popular libraries, we gathered data from three sources for each library: (1) Library Source Code: We collected all available versions of the library source code from GitHub, verifying with PyPI to ensure that the collected versions are formally released and can be installed via pip. From the library source code, we extracted official usage examples for each API from the docstrings. (2) Downstream Application Code: Given Python’s popularity in scientific programming, we collected the source code from top-tier research papers over 10 years as downstream applications. These applications are valuable due to being lightweight yet self-consistent, diverse in their topics, and tagged release timelines associated with publishing venues. Given the time span, this data source implicitly includes evolving libraries. (3) Stack Overflow: Using the library names as queries, we collected FAQ data from Stack Overflow, which provides real user queries and diverse user answers. We filtered the data to include only those queries that explicitly mention the versions of the libraries used, using heuristic rules, as shown in Table 3. Additionally, we have made our best efforts to filter all of the source code based on the open-source licenses of the repositories to ensure there is no infringement.
@@ A.1 Dataset Curation and Collection
Ranked Libraries $\mapsto$ StackOverflow Q&A
@@
Ranked Libraries $\mapsto$ Library Source Code Based on the ranked libraries, parse the source code of these libraries to find functions related to version changes.
Ranked Libraries $\mapsto$ Downstream Application
@@
Table 3: Detailed explanation of annotation stages and the corresponding filtering rules.
[…]
A.2 Lifecycle Tagging of APIs
A.3 Data Preparation for Evaluation
Data Preparation for Token-level Code Completion. As introduced in Section 2, we designed two types of version-controllable code generation tasks: version-specific code completion and version-aware code migration. The task granularities are categorized into token-level, line-level, and block-level to control difficulty and simulate different application scenarios. To better understand model performance, each instance in VersiCode is also tagged with the following: (1) Data source, which includes library source code, downstream applications, and Stack Overflow; (2) Feature type, including addition, deprecation, and general; (3) Release time, i.e. the timestamp from GitHub and Stack Overflow); These tags allow us to filter the evaluation dataset and gain sharper insights into model performance.
[…]
Data Preparation for Code Migration. As shown in Figure 2, considering code migration instances constructed from pairs of metadata, the differences between source and target code versions result in various situations, such as updates from an older version to a newer version or vice versa. Additionally, we categorized versions according to version patterns, for example, treating torch v1.0.0 as a major version and torch v1.3.1 as a minor version, to identify combinations of major and minor version migration cases.
Appendix B Data Statistics and Scope
Dataset Statistics: We present the statistics of VersiCode in Table 4, using the StarCoder2’s (Lozhkov et al., 2024) tokenizer to compute the number of tokens. We also outline the complete version of VersiCode in the table, which furnishes human-labeled data for three additional languages: C#, Java, and JavaScript. Our executable data, applied in Section 4, is a high-quality human-annotated subset from VersiCode, covering 12 libraries, 40 versions, and 119 functionality descriptions. For each functionality description, we matched 4 to 5 test cases.